# Masked Thought: Simply Masking Partial Reasoning Steps Can Improve Mathematical Reasoning Learning of Language Models

### v1

@@ Abstract
In reasoning tasks, even a minor error can cascade into inaccurate results, leading to suboptimal performance of large language models in such domains. Earlier fine-tuning approaches sought to mitigate this by leveraging more precise supervisory signals from human labeling, larger models, or self-sampling, although at a high cost. Conversely, we develop a method that avoids external resources, relying instead on introducing perturbations to the input. Our training approach randomly masks certain tokens within the chain of thought, a technique we found to be particularly effective for reasoning tasks. When applied to fine-tuning with GSM8K, this method achieved a 5% improvement in accuracy over standard supervised fine-tuning with a few codes modified and no additional labeling effort. Furthermore, it is complementary to existing methods. When integrated with related data augmentation methods, it leads to an average improvement of 3% improvement in GSM8K accuracy and 1% improvement in MATH accuracy across five datasets of various quality and size, as well as two base models. We further investigate the mechanisms behind this improvement through case studies and quantitative analysis, suggesting that our approach may provide superior support for the model in capturing long-distance dependencies, especially those related to questions. This enhancement could deepen understanding of premises in questions and prior steps. Our code is available at Github.111 https://github.com/AlibabaResearch/DAMO-ConvAI, and https://github.com/ChangyuChen347/MaskedThought.
@@
[…]
@@ 1 Introduction
To improve the reasoning ability, most training methods focus on acquiring supervised signals obtained through costly means, such as human annotation Liao et al. (2024), generation of larger models Yu et al. (2023); Luo et al. (2023); Liu et al. (2023); Yue et al. (2023), or self-sampling Yuan et al. (2023); Singh et al. (2023); Wang et al. (2023b); Zelikman et al. (2022). These methods depend on these signals to direct learning towards a specific trajectory by precisely outlining the reasoning process. In contrast, our method illustrates that an entirely different approach is equally effective. Instead of requiring more precise guidance to direct the generative process, our technique achieves comparable results by introducing random noises into the input. For reasoning tasks, which are generally considered to require precise understanding, we find that the method of incorporating noise into the reasoning steps is surprisingly effective, indicating that the model might have enhanced its performance through a denoising process. Our approach stands out for its simplicity and effectiveness, offering a feasible alternative to the more complex and resource-intensive methods.
@@
First, we propose a simple and effective Masked thought Fine-Tuning (MFT) method for improving language model reasoning. The implementation of our method is simple: it requires only the substitution of specific tokens with a [mask] in the chain of thought. This is done while maintaining the same procedures as the standard Supervised Fine-Tuning (SFT). With just this minor modification, a 5% increase in accuracy can be achieved when fine-tuning on the GSM8K dataset with Llama-2-7B Touvron et al. (2023), which yields results comparable to those obtained using a more complex data generation pipeline (see Tab.2). Moreover, this method exhibits good versatility and complements other data augmentation techniques well, allowing for seamless integration. Our experiments have demonstrated that by substituting SFT with our Masked Thought strategy, we can enhance accuracy by an average of 3% on the GSM8K dataset and by 1% on the MATH dataset across five different math datasets and two models (see Tab.1). Moreover, MFT demonstrates higher sample efficiency than SFT (see Fig.2).
Second, we analyze our method from a regularization perspective and introduce two guiding principles for effective regularization. Specifically, we propose a noise injection framework in which our method and many commonly employed regularization methods are special cases. These techniques introduce diverse forms of noise to mitigate overfitting, for example, applying dropout to each dimension of the embeddings Srivastava et al. (2014) or employing synonym substitution Zhang et al. (2015). We then conduct a comparative analysis and find that these regularization methods exhibit diminished effectiveness in the absence of the MFT framework (see Tab. 3). Based on our experiments, we propose two guiding principles for an effective regularization: (1) A portion of the tokens in the reasoning path should be retained without noise addition; (2) For positions where noise is added, it is crucial to ensure these positions maintain as less semantics as possible.
@@ 2.1 Masked thought Fine-Tuning
Our approach, named as Masked thought Fine-Tuning, falls under the umbrella of SFT. It maintains the simplicity of standard SFT implementation. To facilitate comparison with other regularization methods, we present a general framework of token noise injection. This framework considers both the position and the type of noise introduced:
@@
$M_{i}\sim$ $\text{Bernoulli}(p)\text{ for each target/source token }w^{gt}_{i}$
Similar to SFT, our method employs the unchanged $w^{gt}$ (the ground truth solotion to question $q$) as the label. The distinction lies in our introduction of noise $n_{i}$ to the input. The implementation of $f$ varies depending on the choice of token noise distribution $\mathcal{N}$ (see Table 3).
This framework is compatible with many commonly used regularization methods. For instance, adding continuous noise to the vectors of all input tokens in both the target and source Jain et al. (2023), where $n_{i}$ follows a Uniform distribution ranging from $-\frac{\alpha}{c}$ to $\frac{\alpha}{c}$, and $M_{i}$ is set to 1. In another approach, scheduled sampling, the model receives its own prediction from the previous time step as input with a probability $p$, here $n_{i}$ represents the predicted token, and $M_{i}$ follows a Bernoulli distribution with parameter $p$ Mihaylova and Martins (2019).
To understand how different choices of $n_{i}$ and $M_{i}$ impact reasoning performance, we conduct a comparative analysis (Tab.  3). Results show that this framework is effective as long as (1) $M_{i}$ should follow a Bernoulli distribution within the target sequence, ensuring that only a subset of tokens in the chain of thought are subjected to noise, and (2) $n_{i}$ is capable of removing the original semantic information at the noisy positions. We find methods that retain some semantics, such as synonym substitution, are less effective.
[…]
@@
(b) The first token is in the mathematical question, and the second is in the chain-of-thought. MFT places greater emphasis on the mathematical question compared to SFT.
Figure 1: We find MFT has a higher long-distance dependency than SFT. A higher bar indicates greater dependency at the corresponding distance. Specifically. We investigate the interdependency of two numerical tokens within a given sequence. We conduct experiments using the Llama-2-7b model trained on the GSM8K dataset with both the SFT and our method.
@@ 2.2 Enhancing Dependency Learning with MFT Regularization
Observation. To understand the underlying reason for the effectivess of MFT, we conduct a detailed analysis and find that models tend to shift towards longer-distance dependencies, particularly enhancing dependency on questions as shown in Fig. 1. Specifically, we iterate over pairs of numeric tokens in a sequence and disturb the first token to see if the second token changes, indicating dependency Schwab and Karlen (2019). As illustrated,  Fig. 1(a) compares the dependency between two tokens within a chain of thought, showing that SFT has a greater dependency on short distances, whereas MFT shifts towards longer-distance dependencies. In  Fig. 1(b), we examine the effect of disturbances in questions on the chain of thought, observing that MFT enhances all dependencies of questions compared to SFT. Then, We check this pattern of other models and datasets in Appendix D. We also observe decreasing trends of short-distance dependencies within the chain. The degree of this decrease differs among various datasets. A clear pattern is that there’s a consistently increased dependence on questions.
@@
Distance Bias. The model may learn this distance bias from the pattern in ground truth data, where previous steps are correctly annotated, allowing the model to extract useful information from nearby thoughts and calculations, sometimes perhaps even overlooking the original question description. However, during inference, it generates steps on its own, moving away from the accurately verified ground truth. With each new step, the likelihood of errors increases. Even without obvious mistakes, inappropriate rephrasing or omissions of question details could mislead subsequent reasoning. This risk is heightened when the model excessively focuses on nearby tokens while neglecting the question, leading to a higher error probability. This phenomenon can also be referred to as exposure bias Bengio et al. (2015), where no safe reasoning is guaranteed during inference.
@@
Dependency Regularization. MFT is helpful in addressing these dependency issues. During step-wise reasoning, an intermediate step might utilize multiple premises from previous steps and questions, forming a directed acyclic graph of dependencies. For the human behavior, in the process of solving mathematical problems, we need to refer both to the initial conditions given in the problem and to many results obtained in previous steps frequently. Although Transformers allow for full connectivity between every token pair, enabling data-driven learning of such dependencies, they risk learning shortcuts through pattern-matching Liu et al. (2022); Khona et al. (2024); Dziri et al. (2023). MFT randomly prunes the fully connected graph during training to mask potential shortcuts, encouraging the model to build up more connections to more robust features, such as information from questions and earlier, less error-prone steps. We maintain the mathematical question unmasked, assuming it is error-free, to prompt the model to comprehend the question while leveraging the provided premises. Empirical evidence shows that masking parts of the questions does not improve performance. However, we find that even though we do not mask the questions, MFT is still able to handle the noise in the questions well (see Tab. 5). For balancing the model’s exploration of unmasked premises and exploitation of the familiar premises, we employ a conservative linear warmup strategy over several epochs for stability. It Allows the model to first see a more complete context, then gradually increase the masking to ensure the stability of the loss value during training.
@@
Table 1: The table presents the results of SFT and our MFT training with various datasets and evaluated on the test sets of GSM8K and MATH. We report the mean across 3 seeds.
@@
Llama2-7B GSM8K 7K - SFT 41.6 -
MFT 47.1 (+5.5) -
MATH 7K - SFT - 7.6
MFT - 8.9 (+1.3)
GSM8K-RFT-U33 110K Llama-7/13/33B SFT 51.9 -
MFT 54.5 (+2.6) -
MAmmoTH 262K GPT-4 SFT 56.2 28.8
@@
Mistral-7B GSM8K 7K - SFT 60.9 -
MFT 64.0 (+3.1) -
MATH 7K - SFT - 16.6
MFT - 17.4 (+0.8)
GSM8K-RFT-U33 110K Llama-7/13/33B SFT 62.9 -
MFT 65.4 (+2.5) -
MAmmoTH 262K GPT-4 SFT 74.9 39.6
@@ 3.1 Dataset
MetaMath: According to Yu et al. (2023), GPT-3.5-Turbo is employed to rephrase queries and responses in both MATH and GSM8K datasets and to introduce reverse engineering problems.
MAmmoTH: Yue et al. (2023) introduces the data named MathInstruct incorporates a variety of math problem datasets, including but not limited to GSM8K, MATH, AQUA Ling et al. (2017), and Camel-Math Li et al. (2023a). It also leverages GPT-4 to generate new reasoning paths through both Program-of-Thought (POT) Chen et al. (2022) and Chain-of-Thought (COT) methods.
Evaluation Dataset. These datasets are mainly augmented from the training sets of GSM8K and MATH, and we test on the corresponding test sets. The GSM8K dataset sets aside 1,319 entries for testing, while the MATH dataset reserves 5,000 entries as a test set. We extract the final answers by specific templates (e.g., “the answer is”) in the predicted reasoning paths following the scripts adopted by Yue et al. (2023); Yu et al. (2023). Subsequently, we calculate the ratio of predicted answers equal to the ground truth answers as the accuracy.
@@ 3.2 Main Results
Improvements Compared with SFT. The results presented in Tab. 1 illustrate that our approach improves performance no matter whether augmentation data is used, demonstrating effectiveness across datasets of varying quality and under two different base models. Notably, this enhancement is most significant in simpler or smaller datasets where the available useful signals in the data may be limited. Our approach demonstrates enhancements in the performance of the MAmmoTH dataset, which incorporates Program-of-Thought Data. This suggests that our method remains effective and is not adversely affected by the inclusion of programming tasks.
@@
Figure 2: Accuracy on GSM8K when training with Mistral-7B and the dataset of MAmmoTH. MFT shows a higher sample efficiency.
@@
Sample Efficiency. Fig 2 illustrates that within an equivalent number of training samples consumed, MFT yields a better performance than SFT. However, this does not imply that our approach requires fewer steps to converge. On the contrary, our method demands additional steps to reach a higher convergence value. This may be attributed to the exploration through masking, enabling us to gather new information from a limited dataset continually. Conversely, extending the training epochs for SFT introduces variability without yielding superior results. In our implementation details, we allow MFT to train for an extended number of epochs to ensure convergence of results as detailed in Tab. C.
@@
Mask Ratio and Scheduling. In Fig. 3, we investigate the influence of two major hyperparameters in MFT: mask ratio and warm-up duration. Our analysis reveals that incorporating a warm-up period contributes to enhanced final performance for a large mask ratio, while a small mask ratio without warm-up also works. Within the GSM8K dataset, varying mask ratios exhibit minimal sensitivity, as the problem-solving procedures within it lack dense formulas. Consequently, this dataset permits a relatively higher mask ratio. In our primary experiment, we set this ratio to 0.4. However, for datasets with more intricate formulas, such as MATH, we opt for smaller mask ratios, typically 0.1 or 0.2. As for warm-up period, we adopt a conservative duration without excessive tuning of this hyperparameter. Generally, we allocate around two-thirds of the total training steps for warm-up.
@@
Figure 3: The impact of mask ratio when training on Llama2-7B with GSM8K. We compare MFT of two settings: Fixed mask ratio without warm-up, and a linear warm-up of the mask ratio starting from 0.
Comparison with Self-Sampling. RFT collects the reasoning paths sampled by the SFT model itself and filters out those paths that contain incorrect answers. Our experiments in Tab. 2 demonstrate that our method performs comparably to the explicit generation of reasoning paths, with even better outcomes when both approaches are applied concurrently. One primary factor of this comparable performance may be that RFT does not sample extensively, yet the critical steps in the reasoning paths remain largely unchanged from the ground truth. The primary distinctions of sampled new paths lie in relatively minor semantic variances, without significantly increasing the new observed connections among local variables Prystawski et al. (2023). It allows us to attain comparable effects by employing masked reasoning paths, where we sample new skip-connection by dropping the intermediate steps. On the other hand, if we compare from the perspective of addressing exposure bias, RL’s self-sampling provides careful supervision for the error-prone inference distribution, whereas MFT aims to directly prevent the model from learning the bias present in the ground truth. We provide additional introduction about RL and RFT in Appendix B.
@@
Despite the performance, for training efficiency, our model is trained on 7K data for 10 epochs, while RFT-100 require training on 46K data for 3 epochs to achieve similar results. Notably, our approach achieves comparable performances without requiring extra machine computing resources.
@@
Table 3: We compared MaskedLM (our main implementation of MFT) with other regularization technique. The probability ($p$) for the sequence noise distribution is fixed at 0.4. The table outlines the framework of Masked thought Fine-Tuning, characterized by: 1. Masking partial tokens during reasoning steps as opposed to introducing noise across all tokens; 2. Prioritizing the minimization of information conveyed by noisy tokens, as indicated by superior outcomes with empty and random strategies.
Sequence Noise Distribution Method Token Noise Distribution $\mathcal{N}$ Hyperparameter GSM8K
@@
$M_{i}=1\quad\text{for each token }i$ Dropout $emb_{i}=emb_{i}\odot D_{i},\quad D_{i}\sim\text{Bernoulli}(q)$ $q$=0.1 40.3
@@
[…]
@@
$M_{i}=1\quad\text{for each target token }i$ Dropout $emb_{i}=emb_{i}\odot D_{i},\quad D_{i}\sim\text{Bernoulli}(q)$ $q$=0.1 39.9
@@
$M_{i}\sim\text{Bernoulli}(p)\text{ for each source token }i,$
where $\text{where }0<p<1$
Dropout $emb_{i}=emb_{i}\odot D_{i},\quad D_{i}\sim\text{Bernoulli}(q)$ $q$=1,$p$=0.2 36.3
@@
;; $M_{i}\sim\text{Bernoulli}(p)\text{ for each target token }i,$
where $\text{where }0<p<1$
[…]
@@
Dropout $emb_{i}=emb_{i}\odot D_{i},\quad D_{i}\sim\text{Bernoulli}(q)$ $q$=0.1 39.0
@@
Scheduled Sampling $seq_{i}=\text{sample}(\text{softmax}(\frac{\text{model output at step }i-1}{\tau}))$ $\tau$=0.1 41.7

### v2

@@ Abstract
In reasoning tasks, even a minor error can cascade into inaccurate results, leading to suboptimal performance of large language models in such domains. Earlier fine-tuning approaches sought to mitigate this by leveraging more precise supervisory signals from human labeling, larger models, or self-sampling, although at a high cost. Conversely, we develop a method that avoids external resources, relying instead on introducing perturbations to the input. Our training approach randomly masks certain tokens within the chain of thought, a technique we found to be particularly effective for reasoning tasks. When applied to fine-tuning with GSM8K on Llama-2-7B, this method achieved a 5% improvement in GSM8K accuracy and a 10% improvement in GSM-IC accuracy over standard supervised fine-tuning with a few codes modified. Furthermore, it is complementary to existing methods. When integrated with related explicit data augmentation methods, it leads to improvements across five datasets of various augmentation methods, as well as two different base models. We further investigate the mechanisms behind this improvement through case studies and quantitative analysis, suggesting that our approach may provide superior support for the model in capturing long-distance dependencies, especially those related to questions. This enhancement could deepen understanding of the premises in questions and prior steps. Our code is available at Github111 https://github.com/AlibabaResearch/DAMO-ConvAI, and https://github.com/ChangyuChen347/MaskedThought..
@@
Changyu Chen1, Xiting Wang1,311footnotemark: 1, Ting-En Lin2, Ang Lv1, Yuchuan Wu2, Xin Gao4, Ji-Rong Wen1, Rui Yan1,3thanks:   Corresponding authors: Xiting Wang, Rui Yan, and Yongbin Li.  and Yongbin Li211footnotemark: 1 1Beijing Key Laboratory of Big Data Management and Analysis Methods, Gaoling School of Artificial Intelligence, Renmin University of China 2Alibaba Group  3Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education 4Computational Bioscience Research Center, KAUST {chen.changyu, xitingwang, anglv,jrwen,ruiyan}@ruc.edu.cn {ting-en.lte, shengxiu.wyc, shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
To improve the reasoning ability, most training methods focus on acquiring supervised signals obtained through costly means, such as human annotation Liao et al. (2024), generation of larger models Yu et al. (2023a); Luo et al. (2023); Liu et al. (2023); Yue et al. (2023), or self-sampling Yuan et al. (2023); Singh et al. (2023); Wang et al. (2023b); Zelikman et al. (2022). These methods enable the models to learn specific trajectories by precisely outlining the reasoning processes. In contrast, our method illustrates that an entirely different approach is equally effective. Instead of requiring more precise guidance to direct the generative process, our technique achieves comparable results by incorporating random noise into the reasoning steps. For reasoning tasks, which are generally considered to require precise understanding, we find that this noise-based method is surprisingly effective, indicating that the model might have enhanced its performance through a denoising process.
@@
First, we propose a simple and effective Masked thought Fine-Tuning (MFT) method for improving language model reasoning. The implementation of our method is simple: it requires only the substitution of specific tokens with a [mask] in the chain of thought. This is done while maintaining the same procedures as the standard Supervised Fine-Tuning (SFT). With this minor modification, MFT produces results comparable to those of a more complex data generation pipeline (see Table 2), and it handles context distractors more effectively (see Table 7). Additionally, our method complements other data augmentation techniques using GPT-3.5 or GPT-4 (see Table 1), and the improvements generalize to out-of-domain reasoning datasets when the instruction set is diverse (see Table 3).
Second, we analyze our method from a regularization perspective and introduce two guiding principles for effective regularization. Specifically, we propose a noise injection framework in which our method and many commonly employed regularization methods are special cases, such as dropout Srivastava et al. (2014) and synonym substitution Zhang et al. (2015). We conduct a comparative analysis and find that these regularization methods are effective only when two guiding principles are met: (1) A portion of the tokens in the reasoning path should be retained without noise addition; (2) For positions where noise is added, it is crucial to ensure these positions maintain as little semantics as possible (see Sec. 3.3).
@@ 2.1 Masked thought Fine-Tuning
Our approach, named Masked thought Fine-Tuning, falls under the umbrella of SFT. It maintains the simplicity of standard SFT implementation. To facilitate comparison with other regularization methods, we present a general framework of token noise injection. This framework considers both the position and the type of noise introduced:
@@
$M_{i}\sim$ $\text{Bernoulli}(p)\text{ for each target/source token }w^{s}_{i}$
Similar to SFT, our method employs the unchanged $w^{gt}$ (the ground truth solution to question $q$) as the label. The distinction lies in our introduction of noise $n_{i}$ to the input. The implementation of $f$ varies depending on the choice of token noise distribution $\mathcal{N}$ (see Tab. 5). We compare various noise settings and ultimately choose [mask] as our method design. The comparisons will be discussed in detail in the experiment section (Sec. 3.3).
@@
(b) The first token is in the mathematical question, and the second is in the chain of thought. MFT places greater emphasis on the mathematical question compared to SFT.
Figure 1: We find MFT has a higher long-distance dependency than SFT. A higher bar indicates greater dependency at the corresponding distance. Specifically. We investigate the dependency of two numerical tokens within a given sequence. We fine-tune Llama-2-7b models on the GSM8K dataset and then count the dependency with 300 samples from the test set.
@@ 2.2 Enhancing Dependency Learning with MFT Regularization
Observation. To understand the underlying reason for the effectiveness of MFT, we conduct a detailed analysis and find that models tend to shift towards longer-distance dependencies, particularly enhancing dependency on questions as shown in Fig. 1. Specifically, we iterate over pairs of numeric tokens in a sequence and disturb the first token to see if the second token changes, indicating dependency Schwab and Karlen (2019). As illustrated,  Fig. 1(a) compares the dependency between two tokens within a chain of thought, showing that SFT has a greater dependency on short distances, whereas MFT shifts towards longer-distance dependencies. In  Fig. 1(b), we examine the effect of disturbances in questions on the chain of thought, observing that MFT enhances all dependencies of questions compared to SFT. Then, We check this pattern of other models and datasets in Appendix D. We also observe decreasing trends of short-distance dependencies within the chain. The degree of this decrease differs among various datasets. A clear pattern is that there’s a consistently increased dependence on questions.
@@
Distance Bias. SFT may cause the model to learn the distance bias from the pattern in ground truth data, where previous steps are correctly annotated, allowing the model to extract useful information from nearby thoughts and calculations, sometimes perhaps even overlooking the original question description. However, during inference, it generates steps on its own, moving away from the accurately verified ground truth. With each new step, the likelihood of errors increases. Even without obvious mistakes, inappropriate rephrasing or omissions of question details could mislead subsequent reasoning. This risk is heightened when the model excessively focuses on nearby tokens while neglecting the question. This phenomenon can also be referred to as exposure bias Bengio et al. (2015), where no safe reasoning is guaranteed during inference.
@@
Dependency Regularization. MFT is helpful in addressing these dependency issues. During step-wise reasoning, an intermediate step might utilize multiple premises from previous steps and questions, forming a directed acyclic graph of dependencies. For human behavior, in the process of solving mathematical problems, we need to refer both to the initial conditions given in the problem and to many results obtained in previous steps frequently. Although Transformers allow for full connectivity between every token pair, enabling data-driven learning of such dependencies, they risk learning shortcuts through pattern-matching Liu et al. (2022); Khona et al. (2024); Dziri et al. (2023). MFT randomly prunes the fully connected graph during training to mask potential shortcuts, encouraging the model to build up more connections to more robust features, such as information from questions and earlier, less error-prone steps. We maintain the mathematical question unmasked, assuming it is error-free, to prompt the model to comprehend the question while leveraging the provided premises. Empirical evidence also shows that masking parts of the questions do not improve performance.
@@
Table 1: The table presents the results of SFT and our MFT training with various datasets and evaluated on the test sets of GSM8K and MATH. ”Augmentation Source“ denotes the model employed to generate data (this can be either new responses or new queries). We report the zero-shot results average across three different seeds.
@@
Llama2-7B GSM8K 7K - SFT 41.6 4.5
MFT 47.1 (+5.5) 4.4 (-0.1)
MATH 7K - SFT 10.1 7.6
MFT 12.9 (+2.8) 8.9 (+1.3)
GSM8K-RFT-U33 110K Llama-7/13/33B SFT 51.9 4.5
MFT 54.5 (+2.6) 4.6 (+0.1)
MathInstruct 262K GPT-4 SFT 56.2 28.8
@@
Mistral-7B GSM8K 7K - SFT 60.9 11.3
MFT 64.0 (+3.1) 10.6 (-0.7)
MATH 7K - SFT 34.7 16.6
MFT 35.8 (+1.1) 17.4 (+0.8)
GSM8K-RFT-U33 110K Llama-7/13/33B SFT 62.9 10.0
MFT 65.4 (+2.5) 10.0 (+0)
MathInstruct 262K GPT-4 SFT 74.9 39.6
@@ 3.1 Datasets and Models
MetaMathQA: According to Yu et al. (2023a), GPT-3.5-Turbo is employed to rephrase queries and responses in both MATH and GSM8K datasets and to introduce reverse engineering problems.
MathInstruct: MAmmoTH Yue et al. (2023) introduces a dataset named MathInstruct, which incorporates a variety of math problem datasets, including but not limited to GSM8K, MATH, AQUA Ling et al. (2017), and Camel-Math Li et al. (2023a). It also leverages GPT-4 to generate reasoning paths using both the Program-of-Thought (POT) Chen et al. (2022b) and Chain-of-Thought (COT) methods.
We also conduct experiments on a non-reasoning dataset in Appendix I.
Evaluation Dataset. These datasets are mainly augmented from the training sets of GSM8K and MATH, and we test on the corresponding test sets. The GSM8K dataset sets aside 1,319 entries for testing, while the MATH dataset reserves 5,000 entries as a test set. Grade-School Math with Irrelevant Context (GSM-IC ) Shi et al. (2023) is a variant of GSM8K testsets containing 58K noisy questions. We extract the final answers by specific templates (e.g., “the answer is”) in the predicted reasoning paths following Yue et al. (2023); Yu et al. (2023a). Subsequently, we calculate accuracy as the ratio of predicted answers matching the ground truth. We test on the Out-of-Domain (OOD) science reasoning dataset following Yue et al. (2024) to check the generalization ability.
Models. We use two types of base models: Llama-2 and Mistral-7B. Llama-2 has a series of widely used base models, and we conduct experiments on its Llama-2-7B  [^2] https://huggingface.co/meta-llama/Llama-2-7b-hf and Llama-2-13B  [^3] https://huggingface.co/meta-llama/Llama-2-13b-hf. Mistral-7B  [^4] https://huggingface.co/mistralai/Mistral-7B-v0.1 is a relatively newer base model with better performance than Llama-2-7B. The default experimental subjects are fine-tuned Llama-2-7B on the GSM8K training set unless otherwise noted in the captions. We pick the checkpoint obtained after a default number of training epochs (3 for SFT following previous works and 10/5 for MFT, as we observe it requires more training steps). For MFT, we default to 10 epochs for training sets smaller than 10K and 5 epochs for other larger datasets. Other hyperparameters are identical for SFT and MFT as listed in implementation details (Appendix C).
@@ 3.2 Main Results
Improvements Compared with SFT. The results presented in Tab. 1 illustrate that our approach improves performance no matter whether augmentation data is used, demonstrating effectiveness across datasets of varying quality and under two different base models. Notably, this enhancement is most significant in simpler or smaller datasets where the available useful signals in the data may be limited. Our approach demonstrates enhancements in the performance of the MathInstruct dataset, which incorporates Program-of-Thought Data. This suggests that our method remains effective when including programming tasks.
@@
Figure 2: Accuracy on GSM8K when training with Mistral-7B and the MathInstruct dataset of MAmmoTH. MFT shows a higher sample efficiency.
Generalization Ability on OOD Reasoning Datasets. To verify whether the improvements brought by our method can generalize to other tasks, we first conduct validations on the single-task training set in Tab.1, i.e., training using GSM8K/MATH and testing on MATH/GSM8K. The experiments show that neither RFT nor MFT contributes to generalization when trained with GSM8K. Additionally, MATH appears to facilitate generalization more effectively compared to GSM8K. We also conduct comparative experiments with multi-source datasets on science reasoning tasks. We select four models trained with multi-source data for comparison. The results in Table 3 indicate that models trained on MathInstruct data show relatively greater improvements compared to those trained on MetaMathQA after MFT training. In contrast, when trained on MetaMathQA with Mistral-7B, there was even a slight decline in performance after MFT training. The difference between these two datasets is that MathInstruct includes 13 sources of mathematical reasoning data, whereas MetaMathQA only includes two datasets. Since MFT is still data-driven, it may require using multi-task data to generalize the knowledge learned from MFT to other reasoning tasks.
Table 3: Test results on OOD reasoning datasets. The testsets include non-mathematical tasks such as biology, physics, and chemistry. Details of each dataset can be found in Appendix L
Model MMLU-S GPQA ARC BBH Avg.
MetaMathQA-Llama2-SFT 33.6 25.2 48.9 32.1 35.0
MetaMathQA-Llama2-MFT 33.4 23.2 51.1 33.4 35.3 (+0.3)
MetaMathQA-Mistral-SFT 53.6 25.7 76.2 50.3 51.4
MetaMathQA-Mistral-MFT 50.7 27.3 72.8 49.6 50.1 (-1.3)
MathInstruct-Llama2-SFT 38.9 22.2 60.6 38.1 40.0
MathInstruct-Llama2-MFT 40.6 25.2 60.0 37.8 40.9 (+0.9)
MathInstruct-Mistral-SFT 48.1 29.8 70.6 47.8 49.1
MathInstruct-Mistral-MFT 52.8 25.7 75.3 52.5 51.6 (+2.5)
@@
Sample Efficiency. Within an equivalent number of training samples consumed, MFT yields a better performance than SFT. However, this does not imply that our approach requires fewer steps to converge. On the contrary, our method demands additional steps to reach a higher convergence value, as illustrated in Fig 2. This may be attributed to the exploration through masking, enabling us to gather new information from a limited dataset continually. Conversely, extending the training epochs for SFT introduces variability without yielding superior results. In our implementation details, we allow MFT to train for an extended number of epochs to ensure convergence of results as detailed in Tab. LABEL:tab:detail.
@@
Mask Ratio and Scheduling. We employ a conservative linear warmup strategy over several epochs for stability. In Fig. 3, we investigate the influence of two major hyperparameters in MFT: mask ratio and warm-up duration. Our analysis reveals that incorporating a warm-up period contributes to enhanced final performance for a large mask ratio, while a small mask ratio without a warm-up also works. Within the GSM8K dataset, varying mask ratios exhibit minimal sensitivity, as the problem-solving procedures within it lack dense formulas. Consequently, this dataset permits a relatively higher mask ratio. In our primary experiment, we set this ratio to 0.4. However, for datasets with more intricate formulas, such as MATH, we opt for smaller mask ratios, typically 0.1 or 0.2. As for the warm-up period, we adopt a conservative duration without excessive tuning of this hyperparameter. Generally, we allocate around two-thirds of the total training steps for warm-up.
@@
Figure 3: The impact of mask ratio when training on Llama2-7B with GSM8K. We compare the MFT of two settings: Fixed mask ratio without warm-up and a linear warm-up of the mask ratio starting from 0.
[…]
@@
Despite the performance, for training efficiency, our model is trained on 7K data for 10 epochs, while RFT-100 requires training on 46K data for 3 epochs to achieve similar results. Notably, our approach achieves comparable performances without requiring extra machine computing resources.
Table 4: Variable Descriptions of compared regularization techniques.
Var. Description
$i$ The index of the token in the sequence.
$M_{i}$ $M_{i}$ is a scalar that equals 1 if noise is injected into the $i$-th token and 0 otherwise.
$p$ $p$ represents a probability, where $\text{where }0<p<1$ and $M_{i}\sim\text{Bernoulli}(p)$.
$emb_{i}$ $emb_{i}$ is the embedding of the $i$-th token.
$D_{i}$ $D_{i}$=$(d_{i}^{1},d_{i}^{2},\ldots,d_{i}^{\lvert emb_{i}\rvert})$ is a vector where, for the $i$-th embedding’s $j$-th dimension, the value is 1 if noise is injected and 0 if not.
$q$ $q$ represents a probability, and $d_{i}^{j}\sim\text{Bernoulli}(q)$.
$\mathcal{U}$ $\mathcal{U}$ represents uniform distribution.
$\alpha$ $\alpha$ is a scalar hyperparameter within the NEFTune framework representing the noise scale.
$c$ $c$ is a scalar as the normalizing factor of NEFTune.
$O_{i}$ $O_{i}$ represents the model output logits over the vocabulary at time step $i$
$\tau$ $\tau$ is a hyperparameter served as the temperature for sampling from the vocabulary distribution. A larger value means larger randomness.
$m$ $m$ is the probability of use [mask].
$V$ $V$ represent all tokens in the vocabulary.
$r$ $r$ is the probability of using a random token in the vocabulary. $m$+$r$=1.
@@
Table 5: We compared MaskedLM (our main implementation of MFT) with other regularization techniques. The probability ($p$) for the sequence noise distribution is fixed at 0.4. The table outlines the framework of Masked thought Fine-Tuning, characterized by: 1. Masking partial tokens during reasoning steps as opposed to introducing noise across all tokens; 2. Prioritizing the minimization of information conveyed by noisy tokens, as indicated by superior outcomes with empty and random strategies. The experiments are conducted on Llama-2-7b with GSM8K.
Sequence Noise Method Token Noise $\mathcal{N}$ Hyperparameter GSM8K
@@
$M_{i}=1\newline \text{for each token }i$ Dropout $emb_{i}=emb_{i}\odot D_{i}$, $D_{i}\sim\text{Bernoulli}(q)$ $q$=0.1 40.3
@@
NEFTune $emb_{i}\mathrel{+}=$ $\mathcal{U}$($-\frac{\alpha}{c},\frac{\alpha}{c})$ $\alpha$=1 40.6
@@
$M_{i}=1\newline \text{for each target token }i$ Dropout $emb_{i}=emb_{i}\odot D_{i}$, $D_{i}\sim\text{Bernoulli}(q)$ $q$=0.1 39.9
@@
$M_{i}\sim\text{Bernoulli}(p)\newline \text{ for each source token }i,$ Dropout $emb_{i}=emb_{i}\odot D_{i}$, $D_{i}\sim\text{Bernoulli}(q)$ $q$=1,$p$=0.2 36.3
@@
$M_{i}\sim\text{Bernoulli}(p)\newline \text{ for each target token }i,$ NEFTune $emb_{i}\mathrel{+}=$ $\mathcal{U}$($-\frac{\alpha}{c},\frac{\alpha}{c})$ $\alpha$=100 42.0
@@
Dropout $emb_{i}=emb_{i}\odot D_{i}$, $D_{i}\sim\text{Bernoulli}(q)$ $q$=0.1 39.0
@@
Scheduled Sampling $seq_{i}\sim\text{softmax}(\frac{O_{i-1}}{\tau})$ $\tau$=0.1 41.7